OpenStudy (anonymous):

factor -16t2 - 32t + 384

OpenStudy (anonymous):

\[ -16 t^2-32 t+384=-16 \left(t^2+2 t-24\right) \]

OpenStudy (anonymous):

\[ -16 t^2-32 t+384=-16 \left(t^2+2 t-24\right)=-16 (t-4) (t+6) \]

OpenStudy (anonymous):

so what is t and what is meaning to the x intercept?

OpenStudy (anonymous):

Nothing is mentioned about x or intercept

OpenStudy (anonymous):

This is the problem that I am trying to solve----A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 384 shows the height f(t), in feet, of the sandbag after t seconds: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function.

OpenStudy (anonymous):

The ball hit the grounds at t=4

OpenStudy (anonymous):

thank you so what does he meaning of the x-intercept of the function.